OpenStudy (anonymous):

plz help me i need to combine like terms but the problem is 2(2x - 4) + x = 7

OpenStudy (anonymous):

4x-8+x=7 5x=7+8 5x=15 x=3

OpenStudy (anonymous):

@shay16

OpenStudy (anonymous):

thx @IQ250